Find the electrostatic force between charges of +2.0C and +5.0 separated by a distance of 75m in a vacuum


1
Expert's answer
2022-02-17T11:36:57-0500

Given:

"q_1=2.0\:\rm C"

"q_2=5.0\:\rm C"

"r=75\:\rm m"


The Coulomb's law says


"F=k\frac{|q_1||q_2|}{r^2}""F=9*10^9\frac{2.0*5.0}{75^2}=1.6*10^7 \:\rm N"

The force is repulsive.